OpenStudy (anonymous):

If b^p=N, then which of the following statements is also true? a. Log (subcript P) N=b b. Log (subscript N) b=p c. N^1/b=p d. Log (subscript b) N=p

OpenStudy (psymon):

Wish there was a good way to, like, explain this without saying the answer xD But here ya go, this is the relationship:

OpenStudy (psymon):

Unless you have a clever way of explaining without just flat out typing the answer @oldrin.bataku

OpenStudy (anonymous):

nah I was just going to hint \(\log b^p=p\log b\)

OpenStudy (psymon):

Fair enough xD

OpenStudy (psymon):

\[\log_{b} N = p\]